# Neural Empirical Interpolation Method for Nonlinear Model Reduction

### v1

@@ 3. Neural empirical interpolation method
Given the setup in the previous section, our goal is to develop a DEIM-inspired neural networks based methodology, namely the Neural Empirical Interpolation Method (NEIM), to find an approximation of the nonlinear term of the form
@@
where $M_{\mu^{(i)}}:\mathbb{R}^{r}\to\mathbb{R}^{r}$ is the neural network corresponding to $\mu^{(i)}$ for $i=1,\dots,k$, and $\theta_{i}(\mu)$ is determined through interpolation of some “optimal” $\theta_{i}$. We reinterpret the affine decomposition assumption from the neural network point of view, through a linear combination of parameter dependent coefficients $\theta_{i}$ and parameter independent networks $M_{\mu^{(i)}}$. Note that $\widehat{\operatorname{N}}(\tilde{v};\mu)\in\mathbb{R}^{r}$, while $\operatorname{N}(U_{r}\tilde{v};\mu)\in\mathbb{R}^{n}$. Since we only work with the truncated POD basis, in what follows, we drop the subscript $r$ in $U_{r}$ for simplicity.
@@ 3.1. NEIM First Step
First, we consider a finite set of parameters =$\mathcal{M}=\{\mu_{i}\}_{i=1}^{m}\subset\mathcal{P}$ to which correspond the snapshots $\mathcal{S}=\{v_{i}=v(\mu_{i})\}_{i=1}^{m}\subset\mathbb{R}^{n}$. We denote their projections onto the POD space by $\tilde{v}_{i}=U^{\top}v_{i}$, $i=1,\dots,m$. When computing the NEIM approximation in the ROM setting, the quantities $\mathcal{M}$ and $\mathcal{S}$ are already available since they are usually needed for the POD approach. Then the first step in NEIM is to initialize our approximation $\widehat{\operatorname{N}}$. We greedily find new terms in the expansion of $\widehat{\operatorname{N}}$, so we denote the iterates with a superscript, the 00-th iterate being $\widehat{\operatorname{N}}^{(0)}\equiv 0$.
@@
(5) $\mu^{(1)}=\operatorname*{\arg\max}_{\mu\in\mathcal{M}}\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\mathcal{E}^{(0)}(v_{i};\mu)^{2},\quad\text{where}\quad\mathcal{E}^{(0)}(v;\mu)=\|U^{\top}\operatorname{N}(v;\mu)-\widehat{\operatorname{N}}^{(0)}(U^{\top}v;\mu)\|_{2}.$
In the above expression, $w_{e}(\mu_{i};\mu)\in[0,\infty)$ are “error” weights which can be interpreted as quadrature weights. In other words, for a fixed $\mu\in\mathcal{M}$, the sum above is an approximation of the $L^{2}$ error over $\mathcal{P}$ between $U^{\top}\operatorname{N}(v(\cdot);\mu)$ and its approximation $\widehat{\operatorname{N}}^{(0)}(U^{\top}v(\cdot);\mu)$, where $v(\cdot)$ denotes the mapping from a parameter $\mu$ to its corresponding high fidelity solution denoted $v(\mu)$. Thus $\mu^{(1)}$ is the parameter such that the corresponding $L^{2}$ error approximation is maximized. Note that by choosing the weights appropriately, we can also obtain a more local error estimator, e.g., by choosing $w_{e}(\mu_{i};\mu)=\delta_{\mu_{i}\mu}$, the Kronecker delta.
Having chosen $\mu^{(1)}$, we train $M_{\mu^{(1)}}$ so that
$M_{\mu^{(1)}}(\tilde{v}_{i})\approx U^{\top}\operatorname{N}(v_{i};\mu^{(1)})/\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2},\quad\text{for}\quad v_{i}\in\mathcal{S}.$
The main reason for evaluating the nonlinearity for a fixed parameter $\mu^{(1)}$ while varying the solution $v_{i}$ is coming from viewing the nonlinear term as an operator; thus the NEIM expansion we obtain approximates the interpolation in $\mu$ of an operator evaluated at particular parameters. To find this first “mode” of the expansion, we take $M_{\mu^{(1)}}$ to be a simple feedforward neural network trained by minimizing the weighted squared loss
@@
At this point we need to find the $\mu$-dependent scaling factor for $M_{\mu^{(1)}}$. Thus, we seek some $\theta_{1}^{(1)}(\mu)$ so that $\theta_{1}^{(1)}(\mu)M_{\mu^{(1)}}(\tilde{v}_{i})$ is “close” to $U^{\top}\operatorname{N}(v_{i};\mu)$ for $\mu\in\mathcal{M}$, and $v_{i}\in\mathcal{S}$. More precisely, we perform an optimization task to find $\theta_{1}^{(1)}(\mu)$ minimizing the following quantity
@@
Then, the new approximation of the projected nonlinearity is $\widehat{\operatorname{N}}^{(1)}(\tilde{v};\mu)=\theta_{1}^{(1)}(\mu)M_{\mu^{(1)}}(\tilde{v})$, $\tilde{v}\in\mathbb{R}^{r}$. Note that $\theta_{1}^{(1)}(\mu)$ is only defined for $\mu\in\mathcal{M}$ because the $\theta$ coefficients are interpolated from $\mathcal{M}$ to all of $\mathcal{P}$ at the end of the NEIM approximation procedure. We can find the value of $\theta_{1}^{(1)}(\mu)$ by setting the partials of (7) equal to 0:
@@
When all of the weights are equal, because we’ve trained $M_{\mu^{(1)}}(\tilde{v}_{i})$ so that $M_{\mu^{(1)}}(\tilde{v}_{i})\approx\frac{U^{\top}\operatorname{N}(v_{i};\mu^{(1)})}{\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2}}$, we expect that $\theta_{1}^{(1)}(\mu^{(1)})\approx\frac{1}{m}\sum_{i=1}^{m}\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2}$. Indeed, taking the weights $w_{e}\equiv C$ gives
@@
Now intuitively, the approximation $\widehat{\operatorname{N}}^{(1)}(\tilde{v};\mu)$ should be accurate for $\mu$ close to $\mu^{(1)}$ and $\tilde{v}$ close to $U^{\top}v$ with $v\in\mathcal{S}$. The first step of NEIM is summarized in Algorithm (2).
@@ 3.2. NEIM Update Step
To add a term to the NEIM expansion, there are two main modifications of the first step described above. First, we must be careful to orthogonalize our training data with respect to the previously trained neural networks in the NEIM expansion to avoid instability in the estimates of the coefficients $\theta_{i}(\mu)$. Second, to solve for the coefficients in the NEIM expansion, we must solve a linear system. Indeed, the first step of NEIM can be viewed as solving a $1\times 1$ linear system for the coefficients. In addition to describing these modifications here, we also generalize the other steps of the NEIM update.
At iteration $j$ of the procedure, we have already computed $\widehat{\operatorname{N}}^{(j-1)}$, and we want to update this approximation by adding a term to the NEIM expansion. We begin, as before, by finding the parameter for which the approximation $\widehat{\operatorname{N}}^{(j-1)}$ has the worst error:
@@
By removing previously selected values for $\mu$ from $\mathcal{M}$ in (8), we avoid selecting again the same parameters, due to e.g. a poorly trained corresponding network, possibly causing stability issues. Moreover, there is no reason why a network with a re-selected parameter would be trained any better than the first network corresponding to that parameter.
Having chosen $\mu^{(j)}$, we want to train $M_{\mu^{(j)}}$ so that
@@
To avoid training instability, $M_{\mu^{(j)}}$ is normalized and orthogonalized with respect to the previously trained neural networks. Specifically, we define $y^{(j)}(\mu_{i})=U^{\top}\operatorname{N}(v_{i};\mu^{(j)})$ for $i=1,\dots,m$, and let $z^{(j)}(\mu_{i})$ be the result of performing Gram-Schmidt on $y^{(j)}(\mu_{i})$ with respect to the networks $M_{\mu^{(1)}}(\tilde{v}_{i}),\dots,M_{\mu^{(j-1)}}(\tilde{v}_{i})$. We describe this in Algorithm (3).
@@
Having orthogonalized the training data, we train $M_{\mu^{(j)}}$ by minimizing the weighted squared loss
@@
where the weights $w_{t}^{(j)}(\mu_{i})\in[0,\infty)$ for $i=1,\dots,m$ have the same meaning as in Equation (6). After training $M_{\mu^{(j)}}$, we find $\theta_{1}^{(j)}(\mu),\dots,\theta_{j}^{(j)}(\mu)$ for $\mu\in\mathcal{M}$ such that
@@
This problem amounts to solving the $j\times j$ linear system resulting from setting the partial derivatives of this expression equal to zero. Doing so, we obtain the following expression for $\mu\in\mathcal{M}$
@@
When the algorithm meets the termination condition given by a suitable s stopping criteria (e.g. the error quadrature (8) is sufficiently small), we either train a neural network or use interpolation to predict the vector of $\theta_{\ell}$ for a newly given parameter $\mu$. The method can be chosen depending on the preferred regression or interpolation context.
The full NEIM training procedure is described in Algorithm (4), while we postpone the discussion regarding the choice of weights and stopping criteria to the next section.
@@
In the NEIM algorithm, we exploit the weighting of the least squares problems in two different steps. Indeed, we weight the training data in the neural network loss at each iteration (Equation (9)), and we also weight the least squares problem for finding the $\theta(\mu)$ coefficients (Equation (10)). Of these, the weighting of the training data is most important practically speaking. To obtain good results, it is needed that the trained neural networks $M_{\mu^{(j)}}$ is accurate when evaluated at $v(\mu^{(j)})$ and $\mu^{(j)}$. Towards this goal, the weight corresponding to $v(\mu^{(j)})$ has to be much larger relative to the other weights.
@@ 3.3. Choice of the Weights and Stopping Criteria
There are many different ways of defining stopping criteria or the most appropriate weighting rules. One could exploit the knowledge of the system and its behavior, or the rule of thumb given in the first remark, but in general this is problem specific (e.g., dependent on the number of snapshots available, the complexity of the solution as parameters vary, etc.). Let us now describe possible ways to choose the weights for the least square procedures. The error weights $w_{e}(\mu_{i};\mu)$ we choose to find the $\theta(\mu)$ coefficients can be interpreted as quadrature weights. We can also interpret training weights $w_{t}^{(j)}(\mu_{i})$ as quadrature weights, but because we want the neural network architectures to be small so that their evaluation is fast in the online phase, it is not necessarily best to use “optimal” quadrature weights. For example, if the parameters $\mu_{i}$ are equi-spaced, then when the neural network is too small and all training points are weighted equally, the network has difficulty fitting any of the data at all. Instead, it may be preferable to give nonzero weights to only a small number of parameters $\mu_{i}$. For example, one could set $w_{t}^{(j)}(\mu_{i}):=\delta_{ij}$ to localize the network training. Another option would be a smooth approximation of the Kronecker delta, for example,
$w_{t}^{(j)}(\mu_{i}):=C\operatorname{exp}\left(-\zeta\|\mu_{i}-\mu_{j}\|_{2}^{2}\right)$
for some scalar parameters $C,\zeta>0$. In the numerical experiments, we specify the weights in each example.
Let us now focus on the different choices to implement stopping criteria for the NEIM algorithm. Within our iterative methods, as we will show in the next section, a measure of the error is provably non-increasing for each subsequent iteration, thus, we can stop NEIM whenever this error is sufficiently small. However, we only know that this error is non-increasing, so it is possible that a desired error tolerance is never reached, so this should be combined with a maximum number of iterations. One could also choose to stop NEIM when this error stops decreasing significantly after each iteration, as in the Elbow Method commonly used in clustering algorithms [37].
@@ 3.4. Error Analysis
We begin by showing that the measure of error we designed for greedily selecting parameters is non-increasing. We then discuss a local error estimate for NEIM to better identify the different contributions to the error in the algorithm.
@@
is non-increasing for each $\mu\in\mathcal{M}$. In particular, if $w_{e}(\mu_{i};\mu)=\delta_{\mu_{i}\mu}$, we have that
$j\mapsto\mathcal{E}^{(j-1)}(v_{i};\mu_{i}),\quad j\in\mathbb{N}$
is non-increasing.
@@
At each iteration of NEIM, for every $\mu\in\mathcal{M}$, we have that $\theta_{1}^{(j)}(\mu),\dots,\theta_{j}^{(j)}(\mu)$ minimize
@@
$\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\mathcal{E}^{(j)}(v_{i};\mu)^{2}$ $=\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\left\|U^{\top}\operatorname{N}(v_{i};\mu)-\sum_{\ell=1}^{j}\theta_{\ell}^{(j)}(\mu)M_{\mu^{(\ell)}}(\tilde{v}_{i})\right\|_{2}^{2}$
$=\operatorname*{\arg\min}_{\theta\in\mathbb{R}^{j}}\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\left\|U^{\top}\operatorname{N}(v_{i};\mu)-\sum_{\ell=1}^{j}\theta_{\ell}M_{\mu^{(\ell)}}(\tilde{v}_{i})\right\|_{2}^{2}$
@@
where in the inequality we took the particular choice of $\theta=(\theta_{1}^{(j-1)},\dots,\theta_{j-1}^{(j-1)},0)\in\mathbb{R}^{j}$. ∎
@@
where $h_{\mu}^{(j)}:\mathbb{R}^{j}\to\mathbb{R}$ has a minimum. In this case, the expression in Equation (8) is replaced by
$\mu^{(j)}=\operatorname*{\arg\max}_{\mu\in\mathcal{M}\setminus\{\mu^{(1)},\dots,\mu^{(j-1)}\}}h_{\mu}^{(j-1)}(\theta_{1}^{(j)}(\mu),\dots,\theta_{j-1}^{(j-1)}(\mu))$
@@
where the $\theta_{\ell}^{(j)}(\mu)$ are to-be-determined by the subsequent steps in the NEIM algorithm.

### v2

@@ 3. Neural empirical interpolation method
Given the setup in the previous section, our goal is to develop a DEIM-inspired neural network based methodology, namely the Neural Empirical Interpolation Method (NEIM), to find an approximation of the nonlinear term of the form
@@
where $M_{\mu^{(i)}}:\mathbb{R}^{r}\to\mathbb{R}^{r}$ is the neural network corresponding to $\mu^{(i)}$ for $i=1,\dots,k$, and $\theta_{i}(\mu)$ is determined through interpolation of some “optimal” $\theta_{i}$. We reinterpret the affine decomposition assumption from the neural network point of view, through a linear combination of parameter dependent coefficients $\theta_{i}$ and parameter independent networks $M_{\mu^{(i)}}$. Note that $\widehat{\operatorname{N}}(\tilde{v};\mu)\in\mathbb{R}^{r}$, while $\operatorname{N}(U_{r}\tilde{v};\mu)\in\mathbb{R}^{n}$. Since we only work with the truncated POD basis, in what follows, we drop the subscript $r$ in $U_{r}$.
@@ 3.1. NEIM First Step
First, we consider a finite set of parameters =$\mathcal{M}=\{\mu_{i}\}_{i=1}^{m}\subset\mathcal{P}$ to which correspond the snapshots $\mathcal{S}=\{v_{i}=v(\mu_{i})\}_{i=1}^{m}\subset\mathbb{R}^{n}$. We denote their projections onto the POD space by $\tilde{v}_{i}=U^{\top}v_{i}$, $i=1,\dots,m$. When computing the NEIM approximation in the ROM setting, the quantities $\mathcal{M}$ and $\mathcal{S}$ are already available from the POD approach. Then the first step in NEIM is to initialize our approximation $\widehat{\operatorname{N}}$. We greedily find new terms in the expansion of $\widehat{\operatorname{N}}$, so we denote the iterates with a superscript, the 00-th iterate being $\widehat{\operatorname{N}}^{(0)}\equiv 0$.
@@
$\mu^{(1)}=\operatorname*{\arg\max}_{\mu\in\mathcal{M}}\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\mathcal{E}^{(0)}(v_{i};\mu)^{2},$
where $\mathcal{E}^{(0)}(v;\mu)=\|U^{\top}\operatorname{N}(v;\mu)-\widehat{\operatorname{N}}^{(0)}(U^{\top}v;\mu)\|_{2}$. In the above expression, $w_{e}(\mu_{i};\mu)\in[0,\infty)$ are “error” weights which can be interpreted as quadrature weights. In other words, for a fixed $\mu\in\mathcal{M}$, the sum above is an approximation of the $L^{2}$ error over $\mathcal{P}$ between $U^{\top}\operatorname{N}(v(\cdot);\mu)$ and its approximation $\widehat{\operatorname{N}}^{(0)}(U^{\top}v(\cdot);\mu)$, where $v(\cdot)$ denotes the mapping from a parameter $\mu$ to its corresponding high fidelity solution denoted $v(\mu)$. Thus $\mu^{(1)}$ is the parameter such that the corresponding $L^{2}$ error approximation is maximized. Note that by choosing the weights appropriately, we can also obtain a more local error estimator, e.g., by choosing $w_{e}(\mu_{i};\mu)=\delta_{\mu_{i}\mu}$, the Kronecker delta. Having chosen $\mu^{(1)}$, we train $M_{\mu^{(1)}}$ so that $M_{\mu^{(1)}}(\tilde{v}_{i})\approx U^{\top}\operatorname{N}(v_{i};\mu^{(1)})/\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2}$ for $v_{i}\in\mathcal{S}$. The main reason for evaluating the nonlinearity for a fixed parameter $\mu^{(1)}$ while varying the solution $v_{i}$ comes from viewing the nonlinear term as an operator; thus the NEIM expansion we obtain approximates the interpolation in $\mu$ of an operator evaluated at particular parameters. To find this first “mode” of the expansion, we take $M_{\mu^{(1)}}$ to be a simple feedforward neural network trained by minimizing the weighted squared loss
@@
At this point we need to find the $\mu$-dependent scaling factor for $M_{\mu^{(1)}}$. Thus, we seek some $\theta_{1}^{(1)}(\mu)$ so that $\theta_{1}^{(1)}(\mu)M_{\mu^{(1)}}(\tilde{v}_{i})$ is “close” to $U^{\top}\operatorname{N}(v_{i};\mu)$ for $\mu\in\mathcal{M}$, and $v_{i}\in\mathcal{S}$. More precisely, we perform an optimization task to find $\theta_{1}^{(1)}(\mu)$ minimizing
@@
The new approximation of the projected nonlinearity is $\widehat{\operatorname{N}}^{(1)}(\tilde{v};\mu)=\theta_{1}^{(1)}(\mu)M_{\mu^{(1)}}(\tilde{v})$, $\tilde{v}\in\mathbb{R}^{r}$. Note that $\theta_{1}^{(1)}(\mu)$ is only defined for $\mu\in\mathcal{M}$ because the $\theta$ coefficients are interpolated from $\mathcal{M}$ to all of $\mathcal{P}$ at the end of the NEIM approximation procedure. We can find the value of $\theta_{1}^{(1)}(\mu)$ by setting the partials of (6) equal to 0:
@@
When all of the weights are equal, having trained $M_{\mu^{(1)}}(\tilde{v}_{i})$ so that $M_{\mu^{(1)}}(\tilde{v}_{i})\approx\frac{U^{\top}\operatorname{N}(v_{i};\mu^{(1)})}{\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2}}$, we expect that $\theta_{1}^{(1)}(\mu^{(1)})\approx\frac{1}{m}\sum_{i=1}^{m}\|U^{\top}\operatorname{N}(v_{i};\mu^{(1)})\|_{2}$. Indeed, taking the weights $w_{e}\equiv C$ gives
@@
Now, the approximation $\widehat{\operatorname{N}}^{(1)}(\tilde{v};\mu)$ should be accurate for $\mu$ close to $\mu^{(1)}$ and $\tilde{v}$ close to $U^{\top}v$ with $v\in\mathcal{S}$. The first step of NEIM is summarized in Algorithm (2).
@@ 3.2. NEIM Update Step
To add a term to the NEIM expansion, there are two main modifications of the first step described above. First, we must be careful to orthogonalize our training data with respect to the previously trained neural networks in the NEIM expansion to avoid instability in the estimates of the coefficients $\theta_{i}(\mu)$. Second, to solve for the coefficients in the NEIM expansion, we must solve a linear system. Indeed, the first step of NEIM can be viewed as solving a $1\times 1$ linear system for the coefficients. In addition to describing these modifications here, we also generalize the other steps of the NEIM update. At iteration $j$ of the procedure, we have already computed $\widehat{\operatorname{N}}^{(j-1)}$, and we want to update this approximation by adding a term to the NEIM expansion. We begin, as before, by finding the parameter for which the approximation $\widehat{\operatorname{N}}^{(j-1)}$ has the worst error:
@@
By removing previously selected values for $\mu$ from $\mathcal{M}$ in (7), we avoid reselecting the same parameters, due to, e.g., a poorly trained corresponding network, possibly causing stability issues. Moreover, there is no reason why a network with a re-selected parameter would be trained any better than the first network corresponding to that parameter. For example, while it is expected that parameters selected by NEIM will correspond to the smallest approximation errors, the results of the neural network training depend on the weights $w_{t}^{(j)}$, and some choices of these weights (e.g., choosing $w_{t}^{(j)}(\mu^{(j)})=0$) can lead to results with the error at the selected parameter not decreasing much after training compared to the error at other parameters. Now, having chosen $\mu^{(j)}$, we want to train $M_{\mu^{(j)}}$ so that
@@
To avoid training instability, $M_{\mu^{(j)}}$ is normalized and orthogonalized with respect to the previously trained neural networks. We define $y^{(j)}(\mu_{i})=U^{\top}\operatorname{N}(v_{i};\mu^{(j)})$ for $i=1,\dots,m$, and let $z^{(j)}(\mu_{i})$ be the result of performing Gram-Schmidt on $y^{(j)}(\mu_{i})$ with respect to the networks $M_{\mu^{(1)}}(\tilde{v}_{i}),\dots,M_{\mu^{(j-1)}}(\tilde{v}_{i})$. We describe this in Algorithm (3).
Remark 1.
Note that the maximum number of terms in the NEIM approximation is $r$ if we assume that the neural networks train well. Indeed, in this case, $\{M_{\mu^{(1)}}(\widetilde{v}_{i}),\dots,M_{\mu^{(r)}}(\widetilde{v}_{i})\}$ would approximate a basis of $\mathbb{R}^{r}$ for each $i=1,\dots,m$, and any additional networks in the NEIM approximation would not greatly reduce the approximation error. This can be compared with DEIM which typically interpolates at many more points than $r$ (with the number of points bounded by the high fidelity dimension). This difference is due to the fact that DEIM approximates a projection of $\operatorname{N}$ in the high fidelity space which is then multiplied by $U^{\top}$, but NEIM directly approximates $U^{\top}\operatorname{N}$ in the reduced order space.
@@
We train $M_{\mu^{(j)}}$ with the orthogonalized data by minimizing the weighted squared loss
@@
with respect to the network parameters, where the weights $w_{t}^{(j)}(\mu_{i})\in[0,\infty)$ for $i=1,\dots,m$ have the same meaning as in Equation (5). After training $M_{\mu^{(j)}}$, we find $\theta_{1}^{(j)}(\mu),\dots,\theta_{j}^{(j)}(\mu)$ for $\mu\in\mathcal{M}$ such that
@@
This problem amounts to solving the $j\times j$ linear system resulting from setting the partial derivatives of this expression equal to zero. Doing so for $\mu\in\mathcal{M}$, we obtain
@@
When the algorithm meets the termination condition given by a suitable stopping criteria (e.g., the error quadrature (7) is sufficiently small), we either train a neural network or use interpolation to predict the vector of $\theta_{\ell}$ for a newly given parameter $\mu$. The method can be chosen depending on the preferred regression or interpolation context. The full NEIM training procedure is described in Algorithm (4). We postpone the discussion regarding weights and stopping criteria to the next section.
@@
In the NEIM algorithm, we exploit the weighting of the least squares problems in two different steps. Indeed, we weight the training data in the neural network loss at each iteration (Equation (8)), and we also weight the least squares problem for finding the $\theta(\mu)$ coefficients (Equation (9)). Of these, the weighting of the training data is most important practically speaking. To obtain good results, it is needed that the trained neural networks $M_{\mu^{(j)}}$ are accurate when evaluated at $\tilde{v}(\mu^{(j)})$ and $\mu^{(j)}$. Towards this goal, the weight corresponding to $\tilde{v}(\mu^{(j)})$ has to be much larger relative to the other weights.
@@ 3.3. Choice of the Weights and Stopping Criteria
There are many different ways of defining stopping criteria or the most appropriate weighting rules. One could exploit the knowledge of the system and its behavior, or the rule of thumb given in the first remark, but in general this is problem specific (e.g., dependent on the number of snapshots available, the complexity of the solution as parameters vary, etc.). Let us now describe possible ways to choose the weights for the least square procedures. The error weights $w_{e}(\mu_{i};\mu)$ we choose to find the $\theta(\mu)$ coefficients can be interpreted as quadrature weights. We can also interpret training weights $w_{t}^{(j)}(\mu_{i})$ as quadrature weights, but because we want the neural network architectures to be small so that their evaluation is fast in the online phase, it is not necessarily best to use “optimal” quadrature weights. For example, if the parameters $\mu_{i}$ are equi-spaced, then when the neural network is too small and all training points are weighted equally, the network has difficulty fitting any of the data at all. Instead, it may be preferable to give nonzero weights to only a small number of parameters $\mu_{i}$. For example, one could set $w_{t}^{(j)}(\mu_{i}):={\delta_{\mu_{i}\mu^{(j)}}}$ to localize the network training. Another option would be a smooth approximation of the Kronecker delta, for some scalar parameters $C,\zeta>0$
$w_{t}^{(j)}(\mu_{i}):=C\operatorname{exp}\left(-\zeta\|\mu_{i}-{\mu^{(j)}}\|_{2}^{2}\right).$
In the numerical experiments, we specify the weights in each example. Let us now focus on the different choices to implement stopping criteria for the NEIM algorithm. Within our iterative methods, as we will show in the next section, a measure of the error is provably non-increasing for each subsequent iteration, thus, we can stop NEIM whenever this error is sufficiently small. However, we only know that this error is non-increasing, so it is possible that a desired error tolerance is never reached, so this should be combined with a maximum number of iterations. One could also choose to stop NEIM when this error stops decreasing significantly after each iteration, as in the Elbow Method commonly used in clustering algorithms [40].
@@ 3.4. Error Analysis
We begin by showing that the measure of error we designed for greedily selecting parameters is non-increasing. We then discuss a local error estimate for NEIM to better identify the different contributions to the error.
@@
is non-increasing for each $\mu\in\mathcal{M}$. In particular, if $w_{e}(\mu_{i};\mu)=\delta_{\mu_{i}\mu}$, we have that $j\mapsto\mathcal{E}^{(j-1)}(v_{i};\mu_{i}),\ j\in\mathbb{N}$ is non-increasing.
@@
At each iteration of NEIM, for every $\mu\in\mathcal{M}$, $\theta_{1}^{(j)}(\mu),\dots,\theta_{j}^{(j)}(\mu)$ minimize
$\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\mathcal{E}^{(j)}(v_{i};\mu)^{2}=\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\left\|U^{\top}\operatorname{N}(v_{i};\mu)-\sum_{\ell=1}^{j}\theta_{\ell}^{(j)}(\mu)M_{\mu^{(\ell)}}(\tilde{v}_{i})\right\|_{2}^{2}.$
@@
$\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)$ $\mathcal{E}^{(j)}(v_{i};\mu)^{2}=\operatorname*{\arg\min}_{\theta\in\mathbb{R}^{j}}\sum_{i=1}^{m}w_{e}(\mu_{i};\mu)\left\|U^{\top}\operatorname{N}(v_{i};\mu)-\sum_{\ell=1}^{j}\theta_{\ell}M_{\mu^{(\ell)}}(\tilde{v}_{i})\right\|_{2}^{2}$
@@
where in the inequality we choose $\theta=(\theta_{1}^{(j-1)},\dots,\theta_{j-1}^{(j-1)},0)\in\mathbb{R}^{j}$. ∎
@@
where $h_{\mu}^{(j)}:\mathbb{R}^{j}\to\mathbb{R}$ has a minimum. Then, the expression in Equation (7) becomes
$\mu^{(j)}=\operatorname*{\arg\max}_{\mu\in\mathcal{M}\setminus\{\mu^{(1)},\dots,\mu^{(j-1)}\}}h_{\mu}^{(j-1)}(\theta_{1}^{{(j-1)}}(\mu),\dots,\theta_{j-1}^{(j-1)}(\mu)),$
@@
where the $\theta_{\ell}^{(j)}(\mu)$ are to-be-determined by the subsequent steps of NEIM.